# Asymptotic safety, quantum gravity, and the swampland: a conceptual assessment

### v2

@@ 3.3 Consequences of no topology change
It is therefore apparent that, while these arguments do not appear to rely on holography, at least directly, they do at least rely on the standard statistical/thermodynamic understanding of semiclassical black-hole physics. Once more, we are led to conclude that, in order to avoid inconsistencies between black-hole thermodynamics and ASQG, one must stipulate that the latter cannot be interpreted in the standard thermodynamic and statistical fashion. In section 4, we shall discuss another independent motivation, grounded in the asymptotic darkness of observables, pointing to the same conclusion.
@@ 3.4 Generalizing ASQG beyond QFT
Is there a generalization of the standard framework(s) of QFT which can get around these issues while still remaining within the framework of ASQG? It seems difficult to answer this question without an extended definition of these terms. Perhaps a seed of this idea is contained in some of the technical approaches used to perform computations in this context, namely those based on the functional RG. In this context, flow equations can define a theory without direct reference to a functional integral over quantum fields. This in principle avoids issues in directly defining functional integrals for gravity, such as the absence of a non-dynamical physical flowing scale. In other words, one may not need coordinates on the space of metrics if one could define observables purely via an auxiliary flow equation. However, this approach still cannot incorporate a sum over topologies directly, [^22] If it could, it would still hardly be related to QFT as we know it, even in its most abstract formulation, because it would lose the functorial assignment of Hilbert spaces, unitary operators and partition functions to manifolds due to the sum over topologies. nor indirectly, since there cannot be any non-field-theoretic degrees of freedom that could replace its effects. In addition, it seems unclear how to encode the physics of extended operators (such as spin holonomies Cheung:2024ypq ) and other global effects (like Dai-Freed anomalies Garcia-Etxebarria:2018ajm ) required by topology fluctuations (see e.g. McNamara:2019rup ; McNamara:2021cuo ) without formulating a more general functional formalism. In a scenario in which such a structure were used to describe something qualitatively different from a functional integral with quantum fields over spacetime, it seems difficult that it would reproduce the semiclassical functional integral which describes the low-energy EFT. However, matrix and tensor models could provide such a framework in which spacetime and topology change emerge along with quantum fields. Specifically, tensor models are also interesting because some may be asymptotically safe Rivasseau:2011hm ; Eichhorn:2013isa ; Eichhorn:2019hsa ; Eichhorn:2020sla ; Gurau:2024nzv ; Eichhorn:2025ezh , and some are related to string theory Banks:1996vh ; Ishibashi:1996xs ; Dijkgraaf:1997vv ; Seiberg:1997ad ; Banks:1999az ; Ydri:2017ncg .
@@ 3.5 Summary and loopholes
The arguments for kinematic swampland constraints that we discussed hinge on spacetime topology change or the thermodynamics of semiclassical black holes, in particular the former implying the latter. Either of these ingredients is crucial to support the absence of global symmetries and the completeness of charge spectra (see figure 2). In the context of ASQG, we argue that the sum over spacetime topologies is very difficult to implement, unless one stretches the definition of QFT, e.g., by including infinitely many fields. Even without such a sum, the considerations based on black-hole entropy, i.e. the asymptotic area law and the existence of microstates (of any possible gauge charge), still apply. In order to circumvent the issues in explaining them with a fundamental QFT of the gravitational field, we can consider a few possible loopholes:
@@ 4 The dynamical swampland
We now turn to dynamical aspects of the swampland and their relation to ASQG. Perhaps as expected, the dynamical aspects of QG are comparatively less clear than the kinematic ones discussed in the preceding section. Important exceptions comprise settings in which gravity is weakly coupled, [^23] Weakly coupled in this context means specifically that the UV cutoff $\Lambda_{\text{EFT}}\ll M_{\text{Pl}}$ be parametrically sub-Planckian. This is because, within EFT, graviton couplings are bounded by $\Lambda_{\text{EFT}}^{2}/M_{\text{Pl}}^{2}\ll 1$. Conversely, a theory with $\Lambda_{\text{EFT}}=M_{\text{Pl}}$ is called strongly coupled. In this precise sense, ASQG is likely strongly coupled. where a number of available results from S-matrix bootstrap methods show an incompatibility between unitary and causal graviton scattering on the one hand, and a strict field theory formulation due to infinite towers of higher-spin species on the other. These ideas are further reinforced by complementary arguments based on black-hole physics Cribiori:2023ffn ; Basile:2023blg ; Basile:2024dqq ; Bedroya:2024ubj ; Herraez:2024kux . Namely, in weakly coupled QG, the smallest black holes in the EFT have size $\Lambda_{\text{EFT}}^{-1}\gg\ell_{\text{Pl}}$, and they transition to (an ensemble of) species below this threshold Herraez:2024kux . Matching thermodynamic quantities Basile:2023blg ; Herraez:2024kux and scattering amplitudes Bedroya:2024ubj at the transition constrains the density of states of these infinite towers of species. These considerations do not lead to an intrinsic issue for the consistency of ASQG, since any such scenario must be non-perturbative anyway. However, they will become important in section 4.2, where weakly coupled limits can appear at infinite distance in theory space. For the time being, we first focus on the role of observables in QG in order to further support the claims made in the preceding section. Then, in section 4.2 we discuss the consequences of information-theoretic factorization in gravity along infinite-distance limits.
@@ 4.1 Observables and UV/IR mixing
What are the observables in QG? No local quantities are gauge invariant, and there are no probes because of the universal coupling to gravity which makes everything interact with everything. [^24] As mentioned in section 3.1, perturbatively one can define observables that are local with respect to a gauge fixing, using e.g. a BRST procedure. Global observables may be (possibly relational) spacetime integrals of local quantities, or they may be holographic, [^25] Here we use the term in a possibly weaker sense than intended, e.g., in McNamaraThesis , where the existence of a local non-gravitational quantum system at the asymptotic boundary is required. i.e. defined purely by data at the asymptotic boundary. In asymptotically AdS superselection sectors, the latter are boundary correlators, as in the AdS/CFT correspondence Witten:1998qj where they arise from a dual conformal field theory; in the asymptotically flat case, they are given by S-matrix elements. It is not clear how this story fits with cosmological de Sitter(-like) settings Banks:2001yp ; Dyson:2002nt ; Dvali:2017eba ; Bedroya:2022tbh ; Antonini:2022fna ; Sahu:2023fbx ; Banks:2024lvl , although perturbatively observable algebras may be defined on observer’s worldlines Witten:2023qsv ; Witten:2023xze .
@@
Asymptotic safety versus asymptotic darkness — In this context, assuming weak (gravitational) coupling at least up to the EFT cutoff, various perturbative results based on the S-matrix bootstrap Camanho:2014apa ; Afkhami-Jeddi:2018apj ; Alonso:2019ptb ; Arkani-Hamed:2020blm ; Geiser:2022icl ; Geiser:2022exp ; Cheung:2022mkw ; Cheung:2023adk ; Cheung:2023uwn ; Haring:2023zwu ; Arkani-Hamed:2023jwn ; Cheung:2024uhn ; Bhardwaj:2024klc point to stringy degrees of freedom: weakly coupled graviton scattering requires infinite towers of higher-spin degrees of freedom Camanho:2014apa ; Caron-Huot:2016icg ; Afkhami-Jeddi:2018apj to preserve causality, while unitarity of the resulting amplitudes seems to only be compatible with a string spectrum and dynamics of the Virasoro-Shapiro type Caron-Huot:2022ugt ; Arkani-Hamed:2020blm ; Geiser:2022exp ; Geiser:2022icl encountered in string theory. Even in the gauge sector, consistency of multiparticle factorization rules out at least some large families of deformations of the Veneziano amplitude which were previously not excluded Arkani-Hamed:2023jwn . However, while existing non-perturbative bounds on Wilson coefficients in the context of high-dimensional maximal supergravity Guerrieri:2021ivu ; Guerrieri:2022sod are compatible with the ranges predicted by string theory, the case for the necessity of non-field-theoretic degrees of freedom is less strong. Since ASQG is a non-perturbative scenario, it may evade the above constraints on its S-matrix. However, due to its field-theoretic origin, the S-matrix of ASQG must exhibit (quantum) scale symmetry in the UV regime, namely it should feature power-law scaling. In particular, this means that (at least) a scattering amplitude should asymptote to a non-vanishing constant or power law at large center-of-mass energy $s\gg\Lambda^{2}_{\text{EFT}}$ (generically $s\gg M^{2}_{\text{Pl}}$einberg:1980gg . This behavior is in clear tension with expectations from asymptotic darkness, also known as black-hole dominance, [^26] The super-exponential degeneracy of intermediate black-hole states implies that the amplitude for any particular final state be suppressed by the the reciprocal of this degeneracy. In particular, a two-to-two amplitude should decay faster than any power law in the (fixed-angle) limit of large center-of-mass energy. namely the idea that the high-energy spectrum of any theory of gravity be dominated by large black holes due to entropy bounds. As a result, in this scenario the high-energy behavior of gravitational amplitudes, at least in certain regimes of impact parameters, would be dominated by black-hole production Banks:1999gd ; Giddings:2007qq ; Giddings:2009gj ; Dvali:2014ila ; Bedroya:2022twb with an effective “single-particle” density of states in $d$-dimensional spacetime scaling as
@@
at high energies, invalidating a field-theoretic integral representation Banks:1999az ; Banks:2010tj . This is because the existence of a Källén-Lehmann integral representation, or a canonical partition function, implies that the single-particle density of states for fields cannot scale super-exponentially. The arguments in Bedroya:2024ubj then translate the same bound to the multiparticle density of states. This argument actually dovetails with our preceding discussion on black-hole entropy in section 3.1, in the sense that both issues would be avoided if ASQG prescribed a different, non-thermodynamic interpretation to these quantities. Notably, such a scaling may be amenable to non-perturbative bootstrap methods in the future, which could further test whether an asymptotically safe field-theoretic S-matrix can be consistent with unitarity and causality principles. As a final remark, in asymptotically AdS sectors the above scaling changes, matching the one of a local conformal field theory on the boundary. However, asymptotic darkness would still make amplitudes decay faster than any power law.
@@ 4.3 Summary and loopholes
In this section we provided arguments to support two main conclusions. The first, which connects to the kinematic considerations in section 3, is that field-theoretic degrees of freedom cannot reproduce the standard super-exponential density of black-hole microstates. In particular, high-energy scattering amplitudes in ASQG must have a scale-invariant high-energy behavior, which is incompatible with black-hole dominance, a.k.a. asymptotic darkness. The second conclusion is that infinite towers of light species cannot arise in all infinite-distance limits, and thus the information-theoretic factorization discussed in Stout:2022phm must be achieved by other means, without infinite towers.
@@
  • outside of weak coupling, addressing black-hole dominance and asymptotic darkness requires loopholes along the lines of section 3.5;
@@ 5 Conclusions
On the swampland side, we considered the network of swampland conjectures (see also figure 2) that are based on bottom-up arguments, i.e., independently of string theory or its patterns. These include the absence of global symmetries, completeness of gauge charges, their relations with black-hole thermodynamics and topology change, and the distance conjecture. The assumption here is that these are universal features of QG in the sense of Eichhorn:2024rkc .
@@
The deep connection between topology change, black-hole thermodynamics, and the “kinematic” swampland conjectures, is schematically summarized in figure 2. Our work thus identifies the first two as root causes of a potential conflict between the above network of swampland conditions and fundamental QFT of gravity. A posteriori, they may be thought of as manifestations of UV/IR mixing expected in QG based on black-hole physics. The comparisons we drew with string theory and AdS/CFT highlight the existence of QG theories which are directly consistent with these swampland conditions. Namely, they are holographic and afford conventional interpretations of black-hole entropy in terms of microstates Strominger:1996sh . A scenario of effective ASQG would fall into this category of QGs as well, since topology fluctuations would be accounted for by the fundamental theory. The corresponding landscape would presumably be compatible with the swampland conjectures we discussed. By contrast, a consistent realization of fundamental ASQG (or similar QFT-based approaches) would constitute a qualitatively different type of QG in these respects. Nevertheless, the landscapes of low-energy EFT could overlap as in figure 1. In summary, either a fundamental description of gravity as a QFT (in the same dimension) is not possible, or the network of swampland constraints considered does not pertain to the universal landscape Eichhorn:2024rkc (or both).
@@ Appendix A A dictionary for non-stringy readers
  • Cobordism: a notion of topological equivalence between manifolds. Two $n$-dimensional compact manifolds $M$, $N$ are said to be cobordant if there exists a manifold $W$ with boundary such that $\partial W=M\sqcup\overline{N}$, where for oriented manifolds the bar denotes the opposite orientation. More precisely, when considering manifolds with some (tangential) structure, such as orientation, (s)pin structure, or a map to some fixed space, one requires that a cobordism $W$ be equipped with the same kind of structure compatible with the restriction to its boundary. Denoting the tangential structure by $\xi$, one writes $M\sim N$ for the equivalence relation of cobordism, and the set of equivalence classes $[M]$ equipped with disjoint union form abelian groups $\Omega^{\xi}_{n}$ called bordism (or cobordism) groups. Bordism groups are useful in physics to encode a number of properties of topological phases of matter Kapustin:2014tfa , anomalies Dai:1994kq as well as topological transitions in QG McNamara:2019rup .

### v3

@@ 3.3 Consequences of no topology change
It is therefore apparent that, while these arguments do not appear to rely on holography, at least directly, they do at least rely on the standard statistical/thermodynamic understanding of semiclassical black-hole physics. Once more, we are led to conclude that, in order to avoid inconsistencies between black-hole thermodynamics and ASQG, one must stipulate that the latter cannot be interpreted in the standard thermodynamic and statistical fashion. In section˜4, we shall discuss another independent motivation, grounded in the asymptotic darkness of observables, pointing to the same conclusion.
@@ 3.4 Generalizing ASQG beyond QFT
Is there a generalization of the standard framework(s) of QFT which can get around these issues while still remaining within the framework of ASQG? It seems difficult to answer this question without an extended definition of these terms. Perhaps a seed of this idea is contained in some of the technical approaches used to perform computations in this context, namely those based on the functional RG. In this context, flow equations can define a theory without direct reference to a functional integral over quantum fields. This in principle avoids issues in directly defining functional integrals for gravity, such as the absence of a non-dynamical physical flowing scale. In other words, one may not need coordinates on the space of metrics if one could define observables purely via an auxiliary flow equation. However, this approach still cannot incorporate a sum over topologies directly, [^23] If it could, it would still hardly be related to QFT as we know it, even in its most abstract formulation, because it would lose the functorial assignment of Hilbert spaces, unitary operators and partition functions to manifolds due to the sum over topologies. nor indirectly, since there cannot be any non-field-theoretic degrees of freedom that could replace its effects. In addition, it seems unclear how to encode the physics of extended operators (such as spin holonomies Cheung:2024ypq ) and other global effects (like Dai-Freed anomalies Garcia-Etxebarria:2018ajm ) required by topology fluctuations (see e.g. McNamara:2019rup ; McNamara:2021cuo ) without formulating a more general functional formalism. In a scenario in which such a structure were used to describe something qualitatively different from a functional integral with quantum fields over spacetime, it seems difficult that it would reproduce the semiclassical functional integral which describes the low-energy EFT. However, matrix and tensor models could provide such a framework in which spacetime and topology change emerge along with quantum fields. Specifically, tensor models are also interesting because some may be asymptotically safe Rivasseau:2011hm ; Eichhorn:2013isa ; Eichhorn:2019hsa ; Eichhorn:2020sla ; Gurau:2024nzv ; Eichhorn:2025ezh , and some others are related to string theory Banks:1996vh ; Ishibashi:1996xs ; Dijkgraaf:1997vv ; Seiberg:1997ad ; Banks:1999az ; Ydri:2017ncg . The former models can be interpreted as a discretization of spacetime, dual to dynamical triangulations; if only the continuum limit is deemed physical, these models fall back into the QFT category. Instead, the latter models are defined on abstract spaces (often of low dimensions, namely zero or one) and spacetime emerges from the collective physics of their degrees of freedom. For instance, the $(0+1)$-dimensional BFSS matrix model Banks:1996vh is conjectured to describe the eleven-dimensional sector of M-theory.
@@ 3.5 Summary and loopholes
The arguments for kinematic swampland constraints that we discussed hinge on spacetime topology change or the thermodynamics of semiclassical black holes, in particular the former implying the latter. Either of these ingredients is crucial to support the absence of global symmetries and the completeness of charge spectra (see figure˜2). In the context of ASQG, we argue that the sum over spacetime topologies is very difficult to implement, unless one stretches the definition of QFT, e.g., by including infinitely many fields. Even without such a sum, the considerations based on black-hole entropy, i.e. the asymptotic area law and the existence of microstates (of any possible gauge charge), still apply. In order to circumvent the issues in explaining them with a fundamental QFT of the gravitational field, we can consider a few possible loopholes:
@@ 4 The dynamical swampland
We now turn to dynamical aspects of the swampland and their relation to ASQG. Perhaps as expected, the dynamical aspects of QG are comparatively less clear than the kinematic ones discussed in the preceding section. Important exceptions comprise settings in which gravity is weakly coupled, [^24] Weakly coupled in this context means specifically that the UV cutoff $\Lambda_{\text{EFT}}\ll M_{\text{Pl}}$ be parametrically sub-Planckian. This is because, within EFT, graviton couplings are bounded by $\Lambda_{\text{EFT}}^{2}/M_{\text{Pl}}^{2}\ll 1$. Conversely, a theory with $\Lambda_{\text{EFT}}=M_{\text{Pl}}$ is called strongly coupled. In this precise sense, ASQG is likely strongly coupled. where a number of available results from S-matrix bootstrap methods show an incompatibility between unitary and causal graviton scattering on the one hand, and a strict field theory formulation due to infinite towers of higher-spin species on the other. These ideas are further reinforced by complementary arguments based on black-hole physics Cribiori:2023ffn ; Basile:2023blg ; Basile:2024dqq ; Bedroya:2024ubj ; Herraez:2024kux . Namely, in weakly coupled QG, the smallest black holes in the EFT have size $\Lambda_{\text{EFT}}^{-1}\gg\ell_{\text{Pl}}$, and they transition to (an ensemble of) species below this threshold Herraez:2024kux . Matching thermodynamic quantities Basile:2023blg ; Herraez:2024kux and scattering amplitudes Bedroya:2024ubj at the transition constrains the density of states of these infinite towers of species. These considerations do not lead to an intrinsic issue for the consistency of ASQG, since any such scenario must be non-perturbative anyway. However, they will become important in section˜4.2, where weakly coupled limits can appear at infinite distance in theory space. For the time being, we first focus on the role of observables in QG in order to further support the claims made in the preceding section. Then, in section˜4.2 we discuss the consequences of information-theoretic factorization in gravity along infinite-distance limits.
@@ 4.1 Observables and UV/IR mixing
What are the observables in QG? No local quantities are gauge invariant, and there are no probes because of the universal coupling to gravity which makes everything interact with everything. [^25] As mentioned in section 3.1, perturbatively one can define observables that are local with respect to a gauge fixing, using e.g. a BRST procedure. Global observables may be (possibly relational) spacetime integrals of local quantities, or they may be holographic, [^26] Here we use the term in a possibly weaker sense than intended, e.g., in McNamaraThesis , where the existence of a local non-gravitational quantum system at the asymptotic boundary is required. i.e. defined purely by data at the asymptotic boundary. These observables can be defined in QFT, in which case they depend on the topology of the spacetime manifold. In asymptotically AdS superselection sectors, the latter are boundary correlators, as in the AdS/CFT correspondence Witten:1998qj where they arise from a dual conformal field theory; in the asymptotically flat case, they are given by S-matrix elements. It is not clear how this story fits with cosmological de Sitter(-like) settings Banks:2001yp ; Dyson:2002nt ; Dvali:2017eba ; Bedroya:2022tbh ; Antonini:2022fna ; Sahu:2023fbx ; Banks:2024lvl , although perturbatively observable algebras may be defined on observer’s worldlines Witten:2023qsv ; Witten:2023xze .
@@
Asymptotic safety versus asymptotic darkness — In this context, assuming weak (gravitational) coupling at least up to the EFT cutoff, various perturbative results based on the S-matrix bootstrap Camanho:2014apa ; Afkhami-Jeddi:2018apj ; Alonso:2019ptb ; Arkani-Hamed:2020blm ; Geiser:2022icl ; Geiser:2022exp ; Cheung:2022mkw ; Cheung:2023adk ; Cheung:2023uwn ; Haring:2023zwu ; Arkani-Hamed:2023jwn ; Cheung:2024uhn ; Bhardwaj:2024klc point to stringy degrees of freedom: weakly coupled graviton scattering requires infinite towers of higher-spin degrees of freedom Camanho:2014apa ; Caron-Huot:2016icg ; Afkhami-Jeddi:2018apj to preserve causality, while unitarity of the resulting amplitudes seems to only be compatible with a string spectrum and dynamics of the Virasoro-Shapiro type Caron-Huot:2022ugt ; Arkani-Hamed:2020blm ; Geiser:2022exp ; Geiser:2022icl encountered in string theory. Even in the gauge sector, consistency of multiparticle factorization rules out at least some large families of deformations of the Veneziano amplitude which were previously not excluded Arkani-Hamed:2023jwn . However, while existing non-perturbative bounds on Wilson coefficients in the context of high-dimensional maximal supergravity Guerrieri:2021ivu ; Guerrieri:2022sod are compatible with the ranges predicted by string theory, the case for the necessity of non-field-theoretic degrees of freedom is less strong. Since ASQG is a non-perturbative scenario, it may evade the above constraints on its S-matrix. However, due to its field-theoretic origin, the S-matrix of ASQG must exhibit (quantum) scale symmetry in the UV regime, namely it should feature power-law scaling. In particular, this means that (at least) a scattering amplitude should asymptote to a non-vanishing constant or power law at large center-of-mass energy $s\gg\Lambda^{2}_{\text{EFT}}$ (generically $s\gg M^{2}_{\text{Pl}}$Weinberg:1980gg . This behavior is in clear tension with expectations from asymptotic darkness, also known as black-hole dominance, [^27] The super-exponential degeneracy of intermediate black-hole states implies that the amplitude for any particular final state be suppressed by the the reciprocal of this degeneracy. In particular, a two-to-two amplitude should decay faster than any power law in the (fixed-angle) limit of large center-of-mass energy. namely the idea that the high-energy spectrum of any theory of gravity be dominated by large black holes due to entropy bounds. As a result, in this scenario the high-energy behavior of gravitational amplitudes, at least in certain regimes of impact parameters, would be dominated by black-hole production Banks:1999gd ; Giddings:2007qq ; Giddings:2009gj ; Dvali:2014ila ; Bedroya:2022twb with an effective “single-particle” density of states in $d$-dimensional spacetime scaling as
@@
at high energies, invalidating a field-theoretic integral representation Banks:1999az ; Banks:2010tj . This is because the existence of a Källén-Lehmann integral representation, or a canonical partition function, implies that the single-particle density of states for fields cannot scale super-exponentially. The arguments in Bedroya:2024ubj then translate the same bound to the multiparticle density of states. This argument actually dovetails with our preceding discussion on black-hole entropy in section˜3.1, in the sense that both issues would be avoided if ASQG prescribed a different, non-thermodynamic interpretation to these quantities. Notably, such a scaling may be amenable to non-perturbative bootstrap methods in the future, which could further test whether an asymptotically safe field-theoretic S-matrix can be consistent with unitarity and causality principles. As a final remark, in asymptotically AdS sectors the above scaling changes, matching the one of a local conformal field theory on the boundary. However, asymptotic darkness would still make amplitudes decay faster than any power law.
@@ 4.3 Summary and loopholes
In this section we provided arguments to support two main conclusions. The first, which connects to the kinematic considerations in section˜3, is that field-theoretic degrees of freedom cannot reproduce the standard super-exponential density of black-hole microstates. In particular, high-energy scattering amplitudes in ASQG must have a scale-invariant high-energy behavior, which is incompatible with black-hole dominance, a.k.a. asymptotic darkness. The second conclusion is that infinite towers of light species cannot arise in all infinite-distance limits, and thus the information-theoretic factorization discussed in Stout:2022phm must be achieved by other means, without infinite towers.
@@
  • outside of weak coupling, addressing black-hole dominance and asymptotic darkness requires loopholes along the lines of section˜3.5;
@@ 5 Conclusions
On the swampland side, we considered the network of swampland conjectures (see also figure˜2) that are based on bottom-up arguments, i.e., independently of string theory or its patterns. These include the absence of global symmetries, completeness of gauge charges, their relations with black-hole thermodynamics and topology change, and the distance conjecture. The assumption here is that these are universal features of QG in the sense of Eichhorn:2024rkc .
@@
The deep connection between topology change, black-hole thermodynamics, and the “kinematic” swampland conjectures, is schematically summarized in figure˜2. Our work thus identifies the first two as root causes of a potential conflict between the above network of swampland conditions and fundamental QFT of gravity. A posteriori, they may be thought of as manifestations of UV/IR mixing expected in QG based on black-hole physics. The comparisons we drew with string theory and AdS/CFT highlight the existence of QG theories which are directly consistent with these swampland conditions. Namely, they are holographic and afford conventional interpretations of black-hole entropy in terms of microstates Strominger:1996sh . A scenario of effective ASQG would fall into this category of QGs as well, since topology fluctuations would be accounted for by the fundamental theory. The corresponding landscape would presumably be compatible with the swampland conjectures we discussed. By contrast, a consistent realization of fundamental ASQG (or similar QFT-based approaches) would constitute a qualitatively different type of QG in these respects. Nevertheless, the landscapes of low-energy EFT could overlap as in figure˜1. In summary, either a fundamental description of gravity as a QFT (in the same dimension) is not possible, or the network of swampland constraints considered does not pertain to the universal landscape Eichhorn:2024rkc (or both).
@@ Appendix A A dictionary for non-stringy readers
  • Cobordism:a notion of topological equivalence between manifolds. Two $n$-dimensional compact manifolds $M$, NN are said to be cobordant if there exists a manifold $W$ with boundary such that $\partial W=M\sqcup\overline{N}$, where for oriented manifolds the bar denotes the opposite orientation. More precisely, when considering manifolds with some (tangential) structure, such as orientation, (s)pin structure, or a map to some fixed space, one requires that a cobordism $W$ be equipped with the same kind of structure compatible with the restriction to its boundary. Denoting the tangential structure by $\xi$, one writes $M\sim N$ for the equivalence relation of cobordism, and the set of equivalence classes $[M]$ equipped with disjoint union form abelian groups $\Omega^{\xi}_{n}$ called bordism (or cobordism) groups. Bordism groups are useful in physics to encode a number of properties of topological phases of matter Kapustin:2014tfa , anomalies Dai:1994kq as well as topological transitions in QG McNamara:2019rup .